OpenStudy (i_ask_alot):

Which formula would you use to determine how much paint would be needed to paint the inside of a circular shape with a diameter of 3 feet on your bedroom wall?

OpenStudy (deathlynine):

We are talking about a circular shape.

OpenStudy (nuttyliaczar):

Basically what is the area of a circle with 3ft diameter?

OpenStudy (deathlynine):

I'm assuming you don't know, and Open Study doesn't like me giving answer. The answer is Pi * r^2.

OpenStudy (nuttyliaczar):

r being half the diameter

OpenStudy (i_ask_alot):

is it C=3^2π

OpenStudy (nuttyliaczar):

The circumference is the length of the curve that makes up the circle. You want the area inside that curve

OpenStudy (mathmale):

It's important that you use the formula given you by another user (above): Pi * r^2. Try again. What is r, the radius of the circle in question?

OpenStudy (i_ask_alot):

I don't know radius of the circle is

OpenStudy (nuttyliaczar):

Diameter is the length from one spot on the edge of the circle to the spot directly across from it. Radius is the distance from the center of the circle to some spot on the edge of the circle. So from that it's easy to see that the radius is just half the diameter

OpenStudy (i_ask_alot):

ok thanks

OpenStudy (i_ask_alot):

here are some answer choices A) C = 2(3)π B) C = 2(1.5)π C) C = 32π D) A = 3/2^2 π

OpenStudy (i_ask_alot):

Sorry C) is 3^2π

OpenStudy (nuttyliaczar):

First off there's only one choice that can even possibly be right since it is solving for the right quantity

OpenStudy (i_ask_alot):

D???

OpenStudy (nuttyliaczar):

Yes because you need the area

OpenStudy (nuttyliaczar):

And see why it works? 3/2 is the diameter cut in half, times pi to complete the formula